OpenStudy (anonymous):

write the following square root in the form a√b, where a and b are integers and b has the least value possible. √20 = ?

jimthompson5910 (jim_thompson5910):

hint: 20 = 4*5 So \[\Large \sqrt{20} = \sqrt{4*5} = \sqrt{4}*\sqrt{5}\]

OpenStudy (anonymous):

i got in my answer \[2\sqrt{5}\]

jimthompson5910 (jim_thompson5910):

you nailed it

OpenStudy (anonymous):

yeah!!!